# Quantifying disorder through sharp spectroscopic fingerprints in an incompressible magnetic state

### v1

@@
Based on the above analysis and exchange parameters, we simulated the FIRMS spectra with and without bond disorder. We employed three different approaches, as summarized in Fig. 3: LSWT, LLD (to account for finite temperature), and ED (to account for quantum corrections). Calculations for a clean system, Fig. 3a-c, show that two sharp single spin-flip modes entirely dominate in the $m\!=\!1/3$ UUD phase, with only weak continua of triple spin-flip excitations in ED at high energies (see SI Sec. IIA). Simulations including disorder paint a different picture, Fig. 3d-f. To account for disorder in LSWT and LLD, we employ a spin vacancy model with two non-adjacent vacancies in a $9\times 9$ supercell. In ED, due to system size constraints, we use a broken-bond model with $d_{zz}=1$ for two bonds in a $6\times 3$ supercell, which equivalently reproduces the same excitations as the vacancy model (see SI Sec. IIC). As the vacancy/disorder density is directly related to the intensity ratio between the dominant and satellite excitations, we change the system size at a fixed number of defects until simulations are in excellent agreement with the data (SI Sec . IIB). Based on this ratio, we find that as little as 2.5% of $\rm Co^{2+}$ vacancies is sufficient to successfully reproduce all aspects of our experimental data in the UUD phase.
@@
We now examine the impact of disorder on thermomagnetic properties of our crystals of KCSO. The isothermal magnetization $M(H)$ measured to sub-Kelvin temperatures with the field along the $c$-axis is shown in Fig. 4a. In the clean case, the magnetization is expected to reach the $m=1/3$ plateau in a single step (see Fig. 1b for a cartoon), as observed by previous experimental studies [26, 27] (see also SI Sec. III). In our case, however, an additional plateau with $m<1/3$ is observed from $\mu_{0}H\approx 2.8$ T, and the main plateau never reaches exactly $m=1/3$. To understand this behavior, we performed classical MC simulations for both our clean and disordered models of KCSO. In the clean case, the magnetization saturates at $m=1/3$ near $\mu_{0}H\approx 4$ T (Fig. 4b). In contrast, for a spin vacancy density of 2.5%, the magnetization shows a plateau around $\mu_{0}H\approx 2.8$ T, followed by a jump to the primary plateau at 3.4 T (Fig. 4c). The magnetization on both plateaus falls below $m=1/3$. By tracking the spin configuration at each magnetic field in our MC simulations, we find that the additional plateau originates from spin flips near vacancy sites (Fig. 4d). This additional plateau serves as a thermodynamic signature of disorder, but lacks the mechanistic insights provided by our spectroscopic approach.
@@
Heat capacity measurements at low temperatures, with the magnetic field applied along the $c$-axis, are presented in Fig. 4f-g. The field dependence of specific heat at $T=0.2$ K reveals that a phase with gapless excitations exists below $\mu_{0}H\approx 1$ T, which we assign to the Y-phase or its quantum analog. As the field increases, the specific heat is suppressed but reaches a finite value again between $\mu_{0}H\approx 2.1$ T and 3.5 T (Fig. 4f). This enhancement survives up to $T\!=\!0.5$ K and matches the peaks observed in the field derivative of magnetization (Fig. 4e). This correlation suggests that the low-temperature specific heat can detect the disorder-induced plateau state through the density of states of local spin-flip processes near vacancies, as these become gapless. Above $T\!=\!1$ K, the heat capacity shows only a broad peak that corresponds to the field-driven transition from the Y phase to the UUD phase. These findings emphasize the significance of examining the low-temperature regime when investigating the effects of disorder, even in systems that exhibit Ising-limit behavior.
@@
The spectacular spectroscopic signatures of disorder we observe in the $m\!=\!1/3$ plateau phase are not unique to KCSO. Additional measurements of the sister compound $\rm Rb_{2}Co(SeO_{3})_{2}$ (See SI Sec. IE) reveal similar satellite modes. Notably, the 2.5% vacancy density extracted in this work appears below the detection limit of laboratory diffraction and microscopy techniques commonly used in the routine structural characterization of quantum materials. By combining the physics of an incompressible magnetization plateau state with the high sensitivity of magneto-optical measurements, we demonstrate a viable route to bridge this technological gap. This approach not only provides quantitative insights into the amount of disorder, but also informs about its mechanistic origin. Given that magnetization plateaus are a common occurrence in frustrated and quantum magnets, and that magneto-optical techniques are becoming more accessible in laboratories worldwide, our method has the potential to accelerate the search for and validation of emergent magnetic states, such as QSLs.
@@ Methods
Single crystals of K2Co(SeO3)2 were prepared following the procedure reported in Ref. [25]. Specifically, high-purity and commercially available K2Co3, Co3O4, and SeO2 were mixed in a molar ratio of $1.2\!:\!0.33\!:\!2.2$ and loaded into an Al2O3 crucible, which was sealed in a vacuumed quartz tube. The tube was heated to $T\!=\!600~{}^{\circ}$, maintained at that temperature for 8 hours, and then cooled to room temperature at a rate of $6~{}^{\circ}$/hour. The resulting flat pink crystals were collected in water. The crystallographic $c^{\ast}$ and $c$ axes are normal to the flat surface of the crystals, as confirmed by backscattering Laue X-ray diffraction. Single crystals of Rb2Co(SeO3)2 were prepared similarly.
@@
Flat samples of appropriate size were selected for far-infrared magneto-spectroscopy (FIRMS) measurements using a Bruker VERTEX 80v Fourier-transform infrared (FTIR) spectrometer at the National High Magnetic Field Laboratory (NHMFL, Tallahassee, Florida). The sample was mounted with GE varnish on a nonmagnetic aperture (1.5 mm in diameter) and attached to a top-loading probe. The probe was then inserted into an optical cryostat equipped with a superconducting magnet reaching up to $\mu_{0}H\!=\!17.5$ T. Infrared light from a mercury lamp was delivered to the sample via an evacuated light pipe and focused onto the sample surface using a parabolic cone.
To reduce light-induced heating and improve signal-to-noise ratio, a 700 cm-1 low-pass filter was placed in the light path just above the sample. Transmitted infrared light passing through the sample was collected by another pair of parabolic cones and detected by a Si bolometer operating at liquid helium temperature. A schematic of the resulting Faraday transmission configuration, with the light propagating along $H\!\parallel\!c$, is shown in Fig. S1. The lowest effective sample temperature during measurements was $T_{\rm min}=5$ K, as measured by a thermometer mounted on the sample holder.
During the measurements, three different beam splitters (Multilayer Mylar, Mylar 50 $\mu$m, and Mylar 125 $\mu$m) were used to cover the relevant energy range. The raw spectra from each (beam splitter) measurement are shown in Fig. S2.
@@
The zero-temperature response ($T\!=\!0$ K) was simulated using linear spin wave theory, by which spin-1/2 operators ${\bf S}_{i}$ are transformed into Holstein-Primakoff bosons ${b}_{i}$ and the resulting bosonic Hamiltonian, truncated at quadratic order, is diagonalized using known numerical approaches implemented in Sunny.jl and other programs [32, 33].
@@
Exact diagonalization calculations were performed using QuSpin.py [30, 31]. The frequency-weighted dynamical spin structure factor, Eq. 5, was calculated using a $6\times 3$ supercell with periodic boundary conditions, employing the Lanczos algorithm [34] and the continued fraction method [35]. The number of Lanczos iterations was set to 200. Due to computational limitations, quenched chemical disorder was modeled using weak bonds rather than directly introducing vacancies. A more detailed discussion can be found in the Supplementary Information Sec. IIC.
@@
A focused ion beam from a Thermo Fisher Helios 5CX was used to lift out a specimen from a K2Co(SeO3)2 single crystal. The specimen was then polished to a thickness of $<50$ nm to achieve atomic resolution using a Hitachi HD-2700 scanning transmission electron microscope (STEM). The microscope was operated at $\Delta V=200$ keV with a convergence angle of 27 mrad and a spatial resolution of $\Delta r\approx 1.3$. Annular dark-field (ADF) images were collected and compared with the K2Co(SeO3)2 lattice structure to determine the zone axis. A selected ADF image is shown in Fig. S10.
@@ Acknowledgements.
We thank N. Peter Armitage for insightful discussions. This work (spectroscopy, thermomagnetic measurements, data analysis, modeling, interpretation and theory) was primarily supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under grants DE-FG02-07ER46451 (SR, NZ, NVS, DS, and ZJ), DE-SC-0018660 (CK and MM), and Early Career Research Program DE-SC-0025478 (AS and IK). The crystal growth effort at UTK (AR, LC, and HZ) was supported by the U.S. Department of Energy grant DE-SC-0020254. The SCXRD measurements at MSU (CP, MX, and WX) were supported the U.S. Department of Energy grant DE-SC-0023648. STEM measurements were performed at the GT Institute for Matter and Systems, a member of the National Nanotechnology Coordinated Infrastructure (NNCI), which is supported by the National Science Foundation under grant NSF-ECCS-2025462. FIRMS measurements were performed at the NHMFL, which is supported by the National Science Foundation Cooperative Agreement under grant NSF-DMR-2128556 and the State of Florida.
@@
SR, NZ, NVS., MO, and DS performed the FIRMS measurements. CK, HZ, and MM carried out the thermodynamic measurements. CK, SR, MM, and ZJ analyzed the experimental data, while CK performed the spin dynamics simulations under the guidance of IK, MM, and ZJ. AS and IK proposed the disorder-bond model. AR, LC, and HZ synthesized the single crystals. CP, MX, and WX performed the SCXRD measurements and fitting. ADV and MT conducted the STEM measurements. CK, SR, MM, and ZJ wrote the manuscript with input from all coauthors. ZJ and DS supervised the project.

### v3

@@ Results
Based on the above analysis and exchange parameters, and the fact that all main and satellite modes in the UUD phase are resolution limited, we simulated the FIRMS spectra with and without vacancy disorder. We employed three different approaches, as summarized in Fig. 4: LSWT, LLD (to account for finite temperature), and ED (to account for quantum corrections). Calculations for a clean system, Fig. 4a-c, show that two sharp single spin-flip modes entirely dominate in the m=1/3m\!=\!1/3 UUD phase, with only weak continua of triple spin-flip excitations in ED at high energies (see Supplementary Sec. IIA). Simulations including disorder paint a different picture, Fig. 4d-f. To account for disorder in LSWT and LLD, we employ a spin vacancy model with non-adjacent vacancies in supercells up to size $18\times 18$. In ED, due to system size constraints, we harness the spectroscopic equivalence between broken bond and vacancy models and use the former with $d_{zz}=1$ for two bonds in a $6\times 3$ supercell, which equivalently reproduces the same excitations as the vacancy model (see Supplementary Sec. IIC). As the vacancy/disorder density is directly related to the intensity ratio between the dominant and satellite excitations, we vary the number of defects (in multiples of 3) in the ($18\times 18$) supercell until the simulations are in excellent agreement with the data (Supplementary Sec. IIB). We find that 2.3(8)% of $\rm Co^{2+}$ vacancies is the most likely range of concentrations, as it successfully reproduces the intensity of all satellite modes in the UUD phase. A more naive analytical approach to estimate the vacancy concentration (Supplementary Sec. IIE), which neglects spectroscopic matrix elements, yields 3.2(14)% of vacancies, within error bars of the more elaborate spin-wave-based method.
@@
We now examine the impact of disorder on the thermomagnetic properties of our crystals of KCSO. Typical thermodynamic measurements are sensitive enough to quantify the impurity concentrations in diluted spin systems [23, 25], and here we extend this concept to a dense spin system. The isothermal magnetization $M(H)$ measured to sub-Kelvin temperatures with the field along the $c$-axis is shown in Fig. 5a. In the clean case, the magnetization is expected to reach the $m=1/3$ plateau in a single step (see Fig. 1b for a cartoon), as observed by previous experimental studies [42, 5] (see also Supplementary Sec. III). In our case, however, an additional plateau with $m<1/3$ is observed from $\mu_{0}H\approx 2.8$ T, and the main plateau never reaches exactly $m=1/3$. To understand this behavior, we performed classical MC simulations for both our clean and disordered models of KCSO with a 2.5% vacancy concentration close to the representative value obtained from FIRMS measurements. In the clean case, the magnetization saturates at $m=1/3$ near $\mu_{0}H\approx 4$ T (Fig. 5b). In contrast, the magnetization in the disordered case shows a pre-plateau around $\mu_{0}H\approx 2.8$ T, followed by a jump to the primary plateau at 3.4 T (Fig. 5c). The magnetization on both plateaus falls below $m=1/3$. By tracking the spin configuration at each magnetic field in our MC simulations, we find that the pre-plateau originates from spin flips near vacancy sites (Fig. 5d). With this understanding, our naive analytical approach (Supplementary Sec. IIE) can be used to estimate the value of the normalized magnetization per spin. In presence of a vacancy concentration $n$, it yields (1n)/3(1-n)/3 and (1n)/32n(1-n)/3-2n for the magnetization at the main plateau and pre-plateau, respectively. This approach best matches our magnetization measurements for $n=2.0\%$ (Fig. 5a), which is remarkably close to the best concentration inferred from FIRMS results, giving further confidence in our vacancy-based disorder model. Thus, the additional plateau we observed serves as a thermodynamic signature of disorder. Taken independently, it lacks the mechanistic insights provided by our spectroscopic approach. But with insight about the disorder mechanism, it yields a surprisingly accurate determination of the disorder concentration.
@@
Heat capacity measurements at low temperatures, with the magnetic field applied along the $c$-axis, are presented in Fig. 5f-g. The field dependence of specific heat at $T=0.2$ K reveals that a phase with gapless excitations exists below $\mu_{0}H\approx 1$ T, which we assign to the Y-phase or its quantum analog. As the field increases, the specific heat is suppressed but reaches a finite value again between $\mu_{0}H\approx 2.1$ T and 3.5 T (Fig. 5f). This enhancement survives up to T=0.5T\!=\!0.5 K and matches the peaks observed in the field derivative of magnetization (Fig. 5e). This correlation suggests that the low-temperature specific heat can detect the disorder-induced plateau state through the density of states of local spin-flip processes near vacancies, as these become gapless. The value of the entropy release between $T\!\approx\!0.2$ and $0.8$ K near $\mu_{0}H\approx 3$ T is consistent with local spin-flip excitations near vacancies (see Supplementary Sec. IV). Above T=1T\!=\!1 K, the heat capacity shows only a broad peak that corresponds to the field-driven transition from the Y phase to the UUD phase. These findings emphasize the significance of examining the low-temperature regime when investigating the effects of disorder, even in systems that exhibit Ising-limit behavior.
Discussion
The spectacular spectroscopic signatures of disorder we observe in the m=1/3m\!=\!1/3 plateau phase are not unique to KCSO. Additional measurements of the sister compound $\rm Rb_{2}Co(SeO_{3})_{2}$ (see Supplementary Sec. IE) reveal similar satellite modes. Notably, the 2.3(8)% vacancy concentration extracted in this work appears below the detection limit of laboratory diffraction and microscopy techniques commonly used in the routine structural characterization of quantum materials. In the UUD phase, this density of vacancies does not jeopardize the original magnetic ground state, unlike other systems where the true ground state of the system may be concealed by disorder effects [26, 11, 10]. Further studies are needed to elucidate if and how the gapless phases of KCSO, such as the V- and Y- phases (and their supersolid counterparts) are affected by the disorder we uncovered.
@@ Discussion
By combining the physics of an incompressible magnetization plateau state with the high sensitivity of magneto-optical measurements, we demonstrate a viable route to probe disorder in magnetic insulators. This approach not only provides quantitative insights into the amount of disorder, but also informs about its mechanistic origin. Given that magnetization plateaus are a common occurrence in frustrated and quantum magnets, and that magneto-optical techniques are becoming more accessible in laboratories worldwide, our method has the potential to accelerate the search for and validation of emergent magnetic states, such as QSLs.
@@ Methods
[…]
@@
FIRMS measurements were performed using a Bruker VERTEX 80v Fourier-transform infrared spectrometer at the National High Magnetic Field Laboratory (NHMFL, Tallahassee, Florida). The sample was mounted with GE varnish on a nonmagnetic aperture (1.5 mm in diameter) and attached to a top-loading probe. The probe was then inserted into an optical cryostat equipped with a superconducting magnet reaching up to $\mu_{0}H\!=\!17.5$ T. Infrared light from a mercury lamp was delivered to the sample via an evacuated light pipe and focused onto the sample surface using a parabolic cone.
To reduce light-induced heating and improve signal-to-noise ratio, a 700 cm-1 low-pass filter was placed in the light path just above the sample. The transmitted infrared light passing through the sample was collected by another pair of parabolic cones and detected by a Si bolometer operating at liquid helium temperature. A schematic of the resulting Faraday transmission configuration, with the light propagating along $H\!\parallel\!c$, is shown in Supplementary Fig. 1. The lowest effective sample temperature during measurements was $T_{\rm min}=5$ K, as measured by a thermometer mounted on the sample holder.
[…]
@@
The zero-temperature response (T=0T\!=\!0 K) was simulated using linear spin wave theory, by which spin-1/2 operators ${\bf S}_{i}$ are transformed into Holstein-Primakoff bosons ${b}_{i}$ and the resulting bosonic Hamiltonian, truncated at quadratic order, is diagonalized using known numerical approaches implemented in Sunny.jl and other programs [30, 24].
@@
Exact diagonalization calculations were performed using QuSpin.py [32, 33]. The frequency-weighted dynamical spin structure factor, Eq. 4, was calculated using a $6\times 3$ supercell with periodic boundary conditions, employing the Lanczos algorithm [18] and the continued fraction method [8]. The number of Lanczos iterations was set to 200. Due to computational limitations, quenched chemical disorder was modeled using weak bonds rather than directly introducing vacancies. A more detailed discussion can be found in the Supplementary Sec. IIC.
@@
A focused ion beam from a Thermo Fisher Helios 5CX was used to lift out a specimen from a K2Co(SeO3)2 single crystal. The specimen was then polished to a thickness of <50<50 nm to achieve atomic resolution using a Hitachi HD-2700 scanning transmission electron microscope (STEM). The microscope was operated at $\Delta V=200$ keV with a convergence angle of 27 mrad and a spatial resolution of $\Delta r\approx 1.3$ Å. Annular dark-field (ADF) images were collected and compared with the K2Co(SeO3)2 lattice structure to determine the zone axis. A selected ADF image is shown in Supplementary Fig. 10.
@@ Acknowledgements.
We thank N. Peter Armitage for insightful discussions. This work (spectroscopy, thermomagnetic measurements, data analysis, modeling, interpretation and theory) was primarily supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under grants DE-FG02-07ER46451 (SR, NZ, NVS, DS, and ZJ), DE-SC0018660 (CK and MM), and Early Career Research Program DE-SC0025478 (theory work by AS and IK). The crystal growth effort at UTK (AR, LC, and HZ) was supported by the U.S. Department of Energy grant DE-SC0020254. The SCXRD measurements at MSU (CP, MX, and WX) were supported the U.S. Department of Energy grant DE-SC0023648. STEM measurements were performed at the GT Institute for Matter and Systems, a member of the National Nanotechnology Coordinated Infrastructure (NNCI), which is supported by the National Science Foundation under grant NSF-ECCS-2025462. FIRMS measurements were performed at the NHMFL, which is supported by the National Science Foundation Cooperative Agreement under grant NSF-DMR-2128556 and the State of Florida.
@@
SR, NZ, NVS., MO, and DS performed the FIRMS measurements. CK, HZ, and MM carried out the thermodynamic measurements. CK, SR, MM, and ZJ analyzed the experimental data, while CK performed the spin dynamics simulations under the guidance of IK, MM, and ZJ. AS and IK proposed and analyzed the disorder-bond model. AR, LC, and HZ synthesized the single crystals. CP, MX, and WX performed the SCXRD measurements and fitting. ADV and MT conducted the STEM measurements. CK, SR, MM, and ZJ wrote the manuscript with input from all coauthors. ZJ and DS supervised the project.